OpenStudy (anonymous):

A dinner plate has a circumference of 106.76 cm. What is the area of the dinner plate? (Use 3.14 for .)

OpenStudy (campbell_st):

well 1st find the radius of the plate using the circumference... \[106.76 = 2 \times \pi \times r\] solve for r. the the area is \[A = \pi r^2\]

OpenStudy (anonymous):

soooo.. 3.14 * 106.76 * 106.76?

OpenStudy (campbell_st):

no \[r = \frac{106.76}{2 \times 3.14}\]